[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jun 28, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jun 28 15:28:53 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. CUMULS locked up - again. Doug Gaer rebooted the workstation.
HSU:
1. GOES-16 data stopped updating in NAWIPS and AWIPS between 01Z and 02Z.  
Note to forecasters: Centers are asked to call SDM ('Tier 1 Support')  
whenever the GOES-16 ISATSS data feed goes down.
Other items:
1. NCO will be performing a bandwidth upgrade on Wednesday, 6/28; network  
service degradation/interruptions are possible during two 20 minute windows  
(1530-1550Z & 1620-1640Z). TAFB/HSU service backup may be required if  
outages last longer than expected. Work was originally to be started on  
Tuesday, June 20 at 1530Z; postponed to Wednesday, June 28 at 1500Z (11 am  
EDT).

2. NCF, NCO, and TSB will be upgrading NHCN to AWIPS build 7.1.3 tomorrow  
starting at 8 AM. Forecasters will need to log out of the workstations at  
that time until the install is complete. Expected completion time: 10 AM.

3. Doug Gaer (NCO) and Chris Mello (TSB) successfully installed AWIPS Build  
17.1.1. on the NHCR testbed system yesterday.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170628/67ad572e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list